In the 2014–2015 academic year, 1,108,165 high school seniors took the SAT. That year the mean of the scores from the math portion was 501 with a standard deviation of 117.

What percentage scored between 600 and 700 points?

*Do NOT include the percentage symbol (%) with your answer.

Round to the second decimal place (0.12).
Answer #1

Solution :

Given that ,

mean = = 501

standard deviation = = 117

P(600 < x < 700) = P[(600 - 501)/ 117) < (x - ) /  < (700 - 501) / 117) ]

= P(0.85 < z < 1.70)

= P(z < 1.70) - P(z < 0.85)

= 0.9554 - 0.8023

= 0.1531

Answer = 0.15